OpenStudy (anonymous):

differentiate 2y=x^4 sin(tan(x^3y))

OpenStudy (anonymous):

2y=x^4 sin(tan(x^3y)) take log both side. log(2y) = 4 sin(tan(x^3y)) *log(x) now differentiate both side w.r.t x d(log(2y))/dx = d(4 sin(tan(x^3y)) *log(x))/dx apply the chain rule of differentiation. (2/y)*y' = 4 cos(tan(x^3y)) *(sec(x^3y))^2 * 3y(x^(3y-1) *y' *log(x) + 4 sin(tan(x^3y)) *1/x we can also write upper eqn in following way. (2/y)*y' - 4 cos(tan(x^3y)) *(sec(x^3y))^2 * 3y(x^(3y-1) *y' *log(x) = 4 sin(tan(x^3y)) *1/x now take y' as common from L.H.S term. and proceed it. y' = 4 sin(tan(x^3y)) *(1/x) /(2/y) - 4 cos(tan(x^3y)) *(sec(x^3y))^2 * 3y(x^(3y-1) *log(x).